Circular No. 6760

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dh IN ANONYMOUS GALAXY O. Perdereau and J. C. Hamilton, on behalf of the Experience de

Recherche d’Objets Sombres (EROS) collaboration, report the discovery on Oct. 20 and 21 CCD images taken by members of the EROS team of an apparent supernova (V ∼ 19), using the Marly 1-m telescope at the European Southern Observatory, La Silla. SN 1997dh is located at α = 23^{h}45^{m}17^{s}.17, δ = −44^{o}56′45′′.1 (equinox 2000.0), which is 4′′ southwest of the center of the host galaxy. Nothing was present at this position on an image (limiting mag ∼ 22) taken on 1996 Oct. 10. A finding chart can be

found at http://www.lal.in2p3.fr/EROS/SNe/Charts/.

RX J0558.0+5353 N. M. Shakhovskoj, Crimean Astrophysical Observatory, Nauchnij;

and S. V. Kolesnikov, Odessa Astronomical Observatory, write: “Periodic variations in the circular polarization, with period P = 0.006301 ± 0.000055 day (544.4 ± 4.8 s) and semiamplitude of 1.80 ± 0.16 percent, were detected in this object on Oct. 4 at the Shajn 2.6-m telescope (+ one-channel polarimeter + fast-rotating achromatic quarter-wave plate and a broad red filter spanning 0.500–0.750 nm). Variations of the circular polarization are sinusoidal with the changing of the polarity. The photometric variations of the same observing run show P = 0.003592 ± 0.000012 day; however, the peaks of the periodogram are accompanied by aliases, one of which corresponds to the double frequency (P = 0.003120 ± 0.000012 day = 269.5 ± 1.0 s) and amplitude 0.035 ± 0.003 mag. This period is very near the period of 272.785 s detected by Ashoka et al. (1995, A.Ap. 297, L83) in white light, and the period of 272.74 s found by Haberl et al. (1994, A.Ap. 291, 171) in x-rays. Light maxima coincide with the extreme positive and negative circular polarization. Johnson magnitudes from Oct. 4 (1.25-m telescope): U = 13.29, B = 13.98, V = 13.91, R = 13.49, I = 13.31.”

SW URSAE MAJORIS J. A. Mattei, AAVSO, reports that this SU UMa-type dwarf nova is

in superoutburst, as indicated by the following visual observations: Oct. 18.51 UT, [14.6 (G. Hanson, Brookfield, WI); 19.17, [13.0 (P. Schmeer, Saarbrücken, Germany); 19.490, 11.1 (Hanson); 19.969, 10.4 (B. Granslo, Nordahl Griegs Vei, Norway); 20.353, 10.5 (Hanson); 22.168, 11.2 (Schmeer, Bischmisheim, Germany). The star reached peak mag 9.6 during its last outburst on 1996 Apr. 13.

1997 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green